One half of a number increased by 16 is four less than two thirds of the number. Find the number. Help please?
I got 1/2x+16=4-2/3x as my equation but I don't know how to solve it.

Answer by nerdybill(7384)

One half of a number increased by 16 is four less than two thirds of the number. Find the number. Help please?
.
Let x = number
then
(1/2)x + 16 = (2/3)x - 4
Multiply both sides by 6:
3x + 96 = 4x - 24
96 = x - 24
120 = x